# Feng-Yu Wanga), Chenggui Yuanb), Xiao-Yu Zhaoa)

### v3

@@ 5.2 Proof of (b)
Let $k>1$. Let $\Theta(t,\cdot):=id+u(t,\cdot)$ for $u$ solving (4.7) such that (4.8) and (4.9) hold. Let
@@
Now, for $\nu_{0}\in\mathscr{P}_{2}(\mathscr{C}_{\tau})$, we choose $\mathscr{F}_{0}$-measurable random variables $X_{0}^{\mu}$ and $X_{0}^{\nu}$ on $\mathscr{C}_{\tau}$ such that
@@
$\leq c_{2}(T)\sup_{t\in[0,T]}\text{\rm{e}}^{-\theta t}\int_{0}^{t}\mathbb{W}_{2}(\mu_{s},\nu_{s})^{2}\,\text{\rm{d}}s\leq\frac{c_{2}(T)}{\theta}\mathbb{W}_{2,\theta}(\mu,\nu)^{2},\ \ \mu,\nu\in\mathcal{C}_{2}^{\gamma}.$
So, when $\theta>c_{2}(T)$, $\Phi^{\gamma}$ is contractive in the complete metric space $(\mathcal{C}_{2}^{\gamma},\mathbb{W}_{2,\theta})$, hence it has a unique fixed point.
@@
$\text{\rm{d}}Y(s)=b(s,Y_{s},\nu_{s})\text{\rm{d}}s+\sigma(s,Y_{s})\text{\rm{d}}\tilde{W}(s),\ \ s\in[0,t],\ \ Y_{0}=X_{0}^{\nu},$

### v5

@@ 5.2 Proof of (b)
Let $k>1$. Let $\Theta(t,\cdot):=id+u(t,\cdot)$ for $u$ solving (4.7) such that (4.8) and (4.9) holds. Let
@@
$\tilde{\mathbb{W}}_{k,\theta,var}(\Phi^{\gamma}\mu,\Phi^{\gamma}\nu)\leq\frac{1}{2}\tilde{\mathbb{W}}_{k,\theta,var}(\mu,\nu),\ \ \mu,\nu\in\mathcal{C}^{\gamma,N}_{k}.$ (5.17)
@@
Now, for $\nu_{0}\in\mathscr{P}_{2}(\mathscr{C}_{\tau})$, for simplicity, we assume that we can choose $\mathscr{F}_{0}$-measurable random variables $X_{0}^{\mu}$ and $X_{0}^{\nu}$ on $\mathscr{C}_{\tau}$ such that
@@
Otherwise, in the following it suffices to first replace $(X_{0}^{\mu},X_{0}^{\nu})$ be the sequences $(X_{0}^{\mu,n},X_{0}^{\nu,n})$ such that
$n^{-1}+\mathbb{W}_{2}(\mu_{0},\nu_{0})^{2}\geq\mathbb{E}\big{[}\|X_{0}^{\mu,n}-X_{0}^{\nu,n}\|_{\tau}^{2}\big{]},\ \ n\geq 1,$
then let $n\rightarrow\infty$.
@@
$\leq c_{2}(T)\sup_{t\in[0,T]}\text{\rm{e}}^{-\theta t}\int_{0}^{t}\mathbb{W}_{2}(\mu_{s},\nu_{s})^{2}\,\text{\rm{d}}s\leq\frac{c_{2}}{\theta}\mathbb{W}_{2,\theta}(\mu,\nu)^{2},\ \ \mu,\nu\in\mathcal{C}_{2}^{\gamma}.$
So, when $\theta>c_{2}$, $\Phi^{\gamma}$ is contractive in the complete metric space $(\mathcal{C}_{2}^{\gamma},\mathbb{W}_{2,\theta})$, hence it has a unique fixed point.
@@
$\text{\rm{d}}Y(s)=b(s,Y_{s},\nu_{s})\text{\rm{d}}s+\sigma(s,Y_{s})\text{\rm{d}}\tilde{W}(s),\ \ s\in[0,t],\ \ Y_{0}=X_{0}^{\nu},$ (6.9)